OpenStudy (anonymous):

7-5t=17 Solve and show work? pls

OpenStudy (anonymous):

7-5t=17 -5t=10 t=-2

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

hey can you solve my other 1?

OpenStudy (anonymous):

sure is it on here already?

OpenStudy (anonymous):

yupp